# Training Machine Learning models at the Edge: A Survey

### v2

@@ VII-C Autonomous vehicles
Autonomous vehicles are vehicles that can operate without human intervention, they utilize sensor technologies, AI, and networking to navigate and make decisions [295]. Autonomous vehicles include self-driving cars, trucks, buses, drones, unmanned aerial vehicles (UAVs), and even small robots. As demonstrated in [296] offloading deep learning tasks to edge devices or servers can improve the inference accuracy while meeting the latency constraint, which makes edge learning prefectly suitable for this use-cases, and as expected there has been extensive research done in this area.
@@
Unmanned Aerial Vehicles (UAVs) are by far the most prominent use of EL for autonomous vehicles. In [297], a Synchronous Federated Learning (SFL) structure for multi-UAVs was proposed, that aims to resolve device privacy concerns that comes from sending raw data to UAV servers, as well as UAVs’ limited processing or communication resources. On the other hand, [298] propose a model-aided federated MARL algorithm to coordinate multiple UAVs on data harvesting missions with limited knowledge about the environment, significantly reducing the real-world training data demand. As mentioned previously in section VII-B, [277] aims to assist farming opperations using UAVs that measure soil moisture and ambient environmental conditions and [299] proposes a model to derive computation specifications for learning-based visual odometry (VO) from physical characteristics of UAVs. Other EL applications for UAVs include [236, 300, 301, 302, 303, 299, 304, 305, 306, 307]
@@
  • Trajectory predictions such as [308] that proposes a solution for trajectory prediction in the edge for both human-driven and autonomous vehicles by leveraging the capabilities of the 5G Multi-Access Edge Computing (5G-MEC) platform to collect and process measurements from vehicles and road infrastructure in edge servers and use a Long Short Term Memory (LSTM) model to predict the vehicle trajectory with high accuracy.
  • Energy efficiency for autonomous vehicles, where [309] proposes a rate-splitting multiple access (RSMA)-based Internet of Vehicles (IoV) system for energy-efficient Federated Edge Learning (FEEL) in autonomous driving, using Non-Orthogonal Unicasting and Multicasting (NOUM) transmission.
VII-D Recommendation systems and personalization
@@
Different approaches using FL have been used for recommendation systems. Amongst them, FedFast [314] propose to accelerate distributed learning for deep federated recommendation models which achieve high accuracy early in the training process. In [315], a Graph Neural Networks (GNNs) used alongside federated learning for social recommendation tasks, a method that aims to alleviate the cold start problem by inducing information of social links between users [315]. On the other hand, [316] was proposed as a federated sequential recommender system for the edge. A method that, unlike traditional recommendations, provides personalized suggestions by sequentially analyzing users’ historical interactions [317]. To achieve this, [316] uses a knowledge-aware transformer and proposed to incorporate knowledge graph information into sequential recommendation tasks, while applying federated learning to preserve users’ privacy, and use replaced token detection and two-stream self-attention strategies to enhance the transformer-based model. Finally, FedCT [123] aims to harness cross-domain recommendation in the edge. While cross-domain recommendation [318] is a promising area for utilizing data from multiple domains, the conventional approach of sharing data between services in a cloud setting often proves impractical or impossible due to privacy and security concerns. This limitation emphasize the appeal of EL as an interesting direction for cross-domain recommendation. By enabling the training of recommender systems on multi-domain data residing on edge devices, while still respecting users’ privacy.
@@
Despite the improvement of federated recommender systems for users’ privacy preservation [319, 320], distributed learning approaches for recommendation still face privacy challenges. Specifically, although users’ item ratings remain on-device, they can be inferred from the final model, thereby posing a risk of data leakage when the model is shared with multiple users [321, 322]. To address this concern, noise is often introduced to the ratings in the form of random user-item interactions. However, this approach usually results in lower performances [322]. In recent years, several solutions have emerged to mitigate this issue. For instance, FedMMF a federated masked matrix factorization, introduced in [321], aims to protect data privacy in federated recommender systems by using personalized mask generated only from local data. Another approach that aims to achieve that is FedRec++ [322], by allocating certain clients as denoising clients to eliminate noise in a privacy-aware manner, thereby counteracting the random sampling of items during the training phase. Other FL based edge recommendation systems includes [320, 323, 319, 324, 325]
@@
Although the most popular approach, FL isn’t the only method to train recommender systems in the edge. In [186], an on-device deep learning sequential recommendation method aimed at mobile devices was proposed, by fine-tuning a pretrained model that was trained using data collected before GDPR  [^2] GDPR : The General Data Protection Regulation is a regulation on data privacy in the European Union and the European Economic Area, for further personalization. And [148] focus on on-device next-item recommendation, and uses compact models and a self-supervised knowledge distillation framework to compensate for the capacity loss caused by compression. Finally, [326] proposes a split-federated learning called SpFedRec where a split learning approach was proposed to migrate the item model from participant’s edge devices to the cloud side and compress item data while transmitting and apply a Squeeze-and-Excitation network mechanism on the backbone model to optimize the perception of dominant features.
@@
Personalized crowdsourced livecast are another part of personalization methods that might benefit from being offloded to the edge. In [327] the rapid development of crowdsourced livecast and the challenges in providing personalized Quality of Experience (QoE) to viewers is discussed, and it introduces an intelligent edge-learning-based framework called ELCast, which integrates Convolutional Neural Networks (CNN) and deep reinforcement learning (DRL) in edge computing architectures for personalized crowdcast recommendation. In the area related to video games, personalisation involves constructing a system capable of adapting video game rules and content to better suit some aspect of the player preferences, personality, experience and performances [328]. Although not yet explored in edge and on-device learning, [329] propose a Deep Q network model to personalize games based on user-interaction on the edge.
VII-E Human Activity Recognition
Human activity recognition (HAR) refers to the automation of the identification and categorization of the various activities performed by humans and their interactions with the environment [330]. Since personalization for HAR has been shown to improve the results and performances of these systems [331], training the model on the edge using Incremental learning or meta learning approaches can help achieve that while increasing the privacy and reducing the bandwidth consumption. PILOTE [127] proposes an incremental learning based approach for HAR that is intended to be trained on edge devices with extremely limited resources and provides reliable performance in handling catastrophic forgetting. While [332] proposes an approach that proposes a personalizable lightweight CNN model for HAR. As well as a training algorithm to find personalization-friendly parameters, for boosting the accuracy after the personalization that is personalized and boost the HAR accuracy when dealing with a wide range of target users. ClusterFL [333], proposes a clustering-based FL approach for HAR training on the edge. Finally, [334] proposes an On-Device deep learning approach for STM32 microcontrollers by fine-tuning a 1-D CNN model for further HAR personalization.
VII-F Others
@@
There are multiple other applications and use cases of edge learning that we couldn’t explore in this section, from Keyword spotting [335, 151], spam detection [40, 41], IoT threats prediction [39], Camera Trap Images classification [336], Detecting Defects in photovoltaic Components [337], Estimating Air Quality [338], to Face Spoof Attack Detection [339] and speech recognition [340, 90, 91, 341, 342, 134, 89], another interesting potential application explored in [343] is the usage of Edge Learning in Lunar analogue environments for future space missions. In general, any use case that benefits from personalization on private user data, or suffers from bandwidth limitations or privacy risks in the training phase might benefit from fully or partially using edge learning. Therefore, we expect the trend of edge learning to continue rising, expend into other fields and areas and grow beyond the current usecases and applications in both academia and the industry.
VIII Libraries, Simulators and Tools for Edge Learning
@@
As an emergent field, Edge Learning (EL) requires multiples tools to facilitate its usability, integration and implementation, ranging from emulators and simulators, used train and test ML models on cloud servers before training on the edge, to libraries that allows the successful training of ML models on edge devices.
VIII-A Libraries and Frameworks for edge learning in edge devices
@@
Although there has been significant work on creating libraries and frameworks for ML at the edge, most of these libraries are primarily focused on the deployment and inference of Deep Learning (DL) on edge devices [344], and only a relative few enable the training of ML models in edge devices. Those tools and libraries include ONNX Runtime [345], TensorFlow Lite [346] or libraries that focus on distributed learning tasks such as flower [347] or FedML [348].
@@
The most popular frameworks for training deep learning models, PyTorch [349] and TensorFlow [350] have both developed libraries for edge learning. However, while TensorFlow lite [346], allows for both the inference and training of DL models at the edge, PyTorch mobile [351] and ExecuTorch [352], Edge ML libraries for PyTorch at the edge, only allow for inference at the time of writing this article. However, it is possible to train PyTorch models using the ONNX Runtime [353]. ONNX Runtime is a cross-platform inference and training machine-learning accelerator, with On-Device Training capabilities [345, 354], the tool has deep integration with PyTorch [353], as well as other deep learning libraries such as Hugging Face [355] and TensorFlow to accelerate training and inferencing on multiple platforms including mobile devices (Android, IOS) and supports multiple hardware accelerators and programming languages.
Federated learning is the most adopted technique to train ML models at the edge (see section V-E), over the years multiple tools and libraries have been proposed to train FL algorithms on the edge. TensorFlow has developed TensorFlow Federated [356] an open-source framework for machine learning and other computations on decentralized data, however, TensorFlow Federated can’t be used to train models on edge devices at the moment of writing this article and is currently intended for experimentation and research uses only. FL_PyTorch [357], a suite of open-source software built on top of PyTorch [349] for Federated Learning, is another library that focus on research, prototyping and experimenting but does not include FL Runtime support [357], other similar libraries include FedJax [358], LEAF [359] and FedLab [360].
Libraries that support the experimentation and development of FL Algorithms, as well running, training and the deployement of these algorithms in edge devices, include the afformentioned ONNX Runtime that allows for FL on edge devices as part of its on-device training support [345, 354]. Flower [347] is a library that is aimed toward research and experimentation of FL algorithms, but allows the execution of the algorithms on a variety of edge devices. FedML [348] is a yet another library that allow researchers to experiment with FL algorithms and migrate them to edge devices. PySyft [361] an FL open-source library built as an extension of some DL libraries, such as PyTorch, Keras and Tensorflow, and can to be run on mobile devices using KotlinSyft  [^3] KotlinSyft: Syft worker for secure on-device machine learning for Android https://github.com/OpenMined/KotlinSyft for Android and SwiftSyft  [^4] SwiftSyft: Syft worker for secure on-device machine learning for iOS https://github.com/OpenMined/SwiftSyft for iOS. And finally, FedERA [362] is a similar library that also includes a verification module for ensuring the validation of local models to avoid aggregating malicious local models and a carbon emission tracker module to accurately estimate CO2 emissions during the local parameter update phase.
Edge learning frameworks and libraries that are aimed at specific platforms or hardware have also been proposed. They include CoreML [363], an ML inference and training framework usable in Swift for iOS, proposed by Apple to simplify the deployment and training of ML models on IOS environment. Aimed at MCUs, PULP-TrainLib was proposed in [364] to enable on-device Training for RISC-V Multi-core MCUs. Librairies and frameworks for approaches other than FL include SplitEasy [101] a framework for training split learning algorithms on mobile devices. And EdgeRL [246] for on-device reinforcement learning, a Light-Weight C/C++ Framework, designed to run on a single-core processor that is typically included in a resource-limited embedded platform.
VIII-B Simulation tools and emulators for edge learning
Emulators and simulation tools are both important in the context of edge computing. simulation tools are used to model the behaviour of fog/edge infrastructures, while emulators provide a way to emulate realistic functionalites and configurations, allowing for the study of interoperability across different layers and protocols in edge-cloud environments [365]. In the context of EL, simulators and emulators, allow for the experimentation of ML models on cloud servers or a single device, while emulating the resources, memory and energy constraints of real edge devices enabling rapid prototyping and experimentation.
@@ VIII-B Simulation tools and emulators for edge learning
Most simulation tools available in the context of edge learning are aimed at federated learning, Flower [347], FedML [348], PySyft [361] and FedERA [362] that have been presented in the section VIII-A,allow running FL algorithms on edge devices in addition to the simulation. FL_PyTorch [357] is a PyTorch [349] based simulation tool for federated learning, and TensorFlow Federated [356] is a similar tool for the TensorFlow framework [350]. Other federated learning simulation tools include : LEAF [359]; FedJax [358] a JAX-based open source library; Flute [366] an open source platform with novel optimization, privacy, and communication strategies ; And finally FedLab [360] a lightweight open-source framework that focus on algorithm effectiveness and communication efficiency, and allows customization on server optimization, client optimization, communication agreement, and communication compression.
IX Open issues, research directions and Future trends
In this section, we dive into the challenges, upcoming research paths, and future trends for Edge Learning (EL). Each part focuses on specific issues, shedding light on the evolving landscape. From handling resource constraints to exploring privacy-friendly applications and combining different techniques, these insights offer a concise look into the forefront of edge learning research.
IX-A Run-time optimization
The resource constraints inherent in edge devices, marked by limited memory, computing capabilities, and often restricted energy availability in mobile edge devices, pose substantial challenges for training ML models. Throughout this paper, we have delved into various techniques and approaches designed to optimize and accelerate the training of ML models on edge devices. However, the focal point of future research around edge learning is anticipated to revolve around the refinement and innovation of these optimization strategies. Given the critical need to mitigate the resource demands of ML training on edge devices, the main research direction is expected to be the development of novel techniques that demand less memory, energy, and computational power. This imperative shift towards more efficient runtime optimization is crucial for ensuring the feasibility and scalability of machine learning on edge devices, ultimately facilitating broader adoption across diverse applications and scenarios.
IX-B Growing Usage of Distributed Methods for Edge Learning
As described in Section V-A, distributed learning methods, such as Federated Learning (FL) and Split Learning (SL), stand out as pivotal techniques in the landscape of edge learning. These methods offer a compelling advantage by distributing the training load of large models across multiple devices, alleviating the computational burden on individual edge devices. Current trends, exemplified in Figure 5, show a steady increase in the adoption of FL and SL in the context of edge learning, with FL emerging as the predominant approach for training ML models at the edge. The ongoing trajectory of this trend is expected to persist, driven by the diverse benefits offered by distributed learning methods. This growing reliance on distributed methodologies not only addresses the challenges posed by resource limitations on edge devices but also paves the way for scalable and efficient ML applications in decentralized environments.
IX-C Hybridization of ML techniques in Edge learning
In the exploration of EL techniques detailed in Section V, various strategies have been employed to optimize the training of relatively large ML models on edge devices. However, each of these techniques, as outlined in Table II, comes with distinct advantages and drawbacks. Recognizing the diversity in these methodologies, there has been a notable surge in approaches that advocate for a hybridization of multiple techniques. Figure 6 illustrates this promising trend, wherein researchers aim to maximize the advantages and mitigate the drawbacks of individual techniques by combining them. This emerging direction, signifies a deliberate effort to create comprehensive and robust solutions tailored to the unique challenges posed by edge devices. Given the promising results showcased by such hybrid approaches, the trajectory indicates a continued surge in interest and research efforts towards refining and expanding the applicability of hybridized techniques in the domain of edge learning.
IX-D Lack of labelled data availability

### v3

@@ V-C Autonomous vehicles
Autonomous vehicles are vehicles that can operate without human intervention, they utilize sensor technologies, AI, and networking to navigate and make decisions [321]. Autonomous vehicles include self-driving cars, trucks, buses, drones, Unmanned Aerial Vehicles (UAVs), and even small robots. As demonstrated in [322] offloading deep learning tasks to edge devices or servers can improve the inference accuracy while meeting the latency constraint, which makes edge learning perfectly suitable for this use-cases, and as expected there has been extensive research done in this area.
@@
UAVs are by far the most prominent use of edge learning for autonomous vehicles. In [323], a synchronous FL structure for multi-UAVs was proposed, that aims to resolve device privacy concerns that come from sending raw data to UAV servers, as well as UAVs’ limited processing or communication resources. On the other hand, [324] propose a model-aided federated MARL algorithm to coordinate multiple UAVs on data harvesting missions with limited knowledge about the environment, significantly reducing the real-world training data demand. As mentioned previously in Section V-B[303] aims to assist farming operations using UAVs that measure soil moisture and ambient environmental conditions and [325] proposes a model to derive computation specifications for learning-based visual odometry from physical characteristics of UAVs. Other edge learning applications for UAVs include [254, 326, 327, 328, 329, 325, 330, 331, 332, 333]
@@
  • Trajectory predictions such as [334] that proposes a solution for trajectory prediction in the edge for both human-driven and autonomous vehicles by leveraging the capabilities of the 5G multi-access edge computing platform to collect and process measurements from vehicles and road infrastructure in edge servers and use an LSTM model to predict the vehicle trajectory with high accuracy.
  • Energy efficiency for autonomous vehicles, where [335] proposes a rate-splitting multiple access (RSMA)-based Internet of Vehicles system for energy-efficient FL in autonomous driving, using non-orthogonal unicasting and multicasting transmission.
V-D Recommendation systems and personalization
@@
Different approaches using FL have been used for recommendation systems. Amongst them, FedFast [340] propose to accelerate distributed learning for deep federated recommendation models which achieve high accuracy early in the training process. In [341], a Graph Neural Networks (GNNs) used alongside FL for social recommendation tasks, a method that aims to alleviate the cold start problem by inducing information of social links between users [341]. On the other hand, [342] was proposed as a federated sequential recommender system for the edge. A method that, unlike traditional recommendations, provides personalized suggestions by sequentially analyzing users’ historical interactions [343]. To achieve this, [342] uses a knowledge-aware transformer and proposed to incorporate knowledge graph information into sequential recommendation tasks, while applying FL to preserve users’ privacy, and use replaced token detection and two-stream self-attention strategies to enhance the transformer-based model. Finally, FedCT [133] aims to harness cross-domain recommendation in the edge. While cross-domain recommendation [344] is a promising area for utilizing data from multiple domains, the conventional approach of sharing data between services in a cloud setting often proves impractical or impossible due to privacy and security concerns. This limitation emphasizes the appeal of edge learning as an interesting direction for cross-domain recommendation. By enabling the training of recommender systems on multi-domain data residing on edge devices, while still respecting users’ privacy.
@@
Despite the improvement of federated recommender systems for users’ privacy preservation [345, 346], distributed learning approaches for recommendation still face privacy challenges. Specifically, although users’ item ratings remain on-device, they can be inferred from the final model, thereby posing a risk of data leakage when the model is shared with multiple users [347, 348]. To address this concern, noise is often introduced to the ratings in the form of random user-item interactions. However, this approach usually results in lower performances [348]. In recent years, several solutions have emerged to mitigate this issue. For instance, FedMMF a federated masked matrix factorization, introduced in [347], aims to protect data privacy in federated recommender systems by using personalized mask generated only from local data. Another approach that aims to achieve that is FedRec++ [348], by allocating certain clients as denoising clients to eliminate noise while respecting privacy, thereby counteracting the random sampling of items during the training phase. Other FL based edge recommendation systems includes [346, 349, 345, 350, 351]
@@
Although the most popular approach, FL isn’t the only method to train recommender systems in the edge. In [198], an on-device deep learning sequential recommendation method aimed at mobile devices was proposed, by fine-tuning a pretrained model that was trained using data collected before GDPR [^3] GDPR: The General Data Protection Regulation is a regulation on data privacy in the European Union and the European Economic Area, for further personalization. And [160] focus on on-device next-item recommendation, and uses compact models and a self-supervised knowledge distillation framework to compensate for the capacity loss caused by compression. Finally, [352] proposes a split-FL method called SpFedRec where a split learning approach was proposed to migrate the item model from participant’s edge devices to the cloud side and compress item data while transmitting and apply a Squeeze-and-Excitation network mechanism on the backbone model to optimize the perception of dominant features.
@@
Personalized crowdsourced livecast are another part of personalization methods that might benefit from being offloaded to the edge. In [353] the rapid development of crowdsourced livecast and the challenges in providing personalized quality of experience to viewers is discussed, and it introduces an intelligent edge-learning-based framework called ELCast, which integrates CNNs and deep RL models in edge computing architectures for personalized crowdcast recommendation. In the area related to video games, personalization involves constructing a system capable of adapting video game rules and content to better suit some aspect of the player preferences, personality, experience and performances [354]. Although not yet explored in edge and on-device learning, [355] propose a Deep Q network model to personalize games based on user-interaction on the edge.
V-E Others
@@
There are multiple other applications and use cases of edge learning that we couldn’t explore in this section, from keyword spotting [356, 163], spam detection [56, 57], IoT threats prediction [55], camera trap images classification [357], detecting defects in photovoltaic components [358], estimating air quality [359], to face spoof attack detection [360] and speech recognition [361, 94, 95, 362, 363, 144, 93], another interesting potential application explored in [364] is the use of edge learning in lunar analogue environments for future space missions. In general, any use case that benefits from personalization on private user data, or suffers from bandwidth limitations or privacy risks in the training might benefit from fully or partially using edge learning. Therefore, we expect the trend of edge learning to continue rising, expend into other fields and areas and grow beyond the current use cases and applications in both academia and the industry.
VI Libraries, Simulators and Tools for Edge Learning
@@
As an emergent field, edge learning requires multiples tools to facilitate its usability, integration and implementation, ranging from emulators and simulators, used train and test ML models on cloud servers before training on the edge, to libraries that allows the successful training of ML models on edge devices.
@@
Although there has been significant work on creating libraries and frameworks for ML at the edge, most of these libraries focus on the deployment and inference of deep learning on edge devices [365]. Only a few libraries enable the training of ML models on edge devices. These include ONNX Runtime [^4] https://cloudblogs.microsoft.com/opensource/2023/05/31/on-device-training-efficient-training-on-the-edge-with-onnx-runtime/, TensorFlow Lite [^5] https://blog.tensorflow.org/2021/11/on-device-training-in-tensorflow-lite.html or libraries that focus on distributed learning tasks such as Flower [366] or FedML [367]. Some tools, only allow for researching, prototyping and experimenting of FL methods and are designed for simulating FL methods on the cloud. While others allow for the training and deployment of these techniques in edge devices. Table II shows a list of frameworks intended for edge learning, or for running training simulations for edge learning.
TABLE III: Comparison between The different Frameworks for edge learning
Framework Name Support simulation Allow training on edge device Type Language Plateform
TensorFlow lite Deep Learning Python Android / iOS
ONNX Runtime Deep Learning / ML Multiple Languages Android / iOS / Other Plateforms
Flower [366] FL Python Android / iOS
FedML [367] FL Python Android / iOS
PySyft [368] FL Python, Kotlin, Swift Android / iOS
FedERA [369] FL Python, Kotlin, Swift Android / iOS
FedLab [370] FL Python -
FedJax [371] FL Python, Jax -
LEAF [372] FL Python -
Flute [373] FL Python -
PyVertical [374] FL / Split Learning Python -
OpenFL [375] FL Python -
EasyFL [376] FL Python -
FL_PyTorch [377] FL Python -
TensorFlow Federated FL Python -
CoreML ML Python iOS
EdgeRL [378] RL C/C++ Embedded Platforms
PULP-TrainLib [379] Deep Learning C RISC-V Multi-core MCUs
@@
PyTorch [380] and TensorFlow [381], the most popular frameworks for training deep learning models, have both developed edge ML libraries. However, while TensorFlow lite, allows for both the inference and training on the edge, PyTorch mobile [^6] https://pytorch.org/mobile/home/ and ExecuTorch [^7] https://pytorch.org/blog/pytorch-edge/, Edge ML libraries for PyTorch at the edge, only support inference at the time of writing. Note that PyTorch models can be trained using ONNX Runtime [^8] https://onnxruntime.ai/blogs/pytorch-on-the-edge. ONNX Runtime is a cross-platform ML accelerator with on-device training capabilities. It has deep integration with PyTorch, Hugging Face [^9] https://huggingface.co/blog/optimum-onnxruntime-training and TensorFlow, enabling accelerated training and inference on multiple platforms, including mobile devices (Android, iOS) and various hardware accelerators and programming languages. Additionally, ONNX Runtime supports FL on edge devices through its on-device training capabilities.
Over the years, multiple tools and libraries have been proposed to train FL algorithms on edge devices, driven by the need for efficient and decentralized learning. As mentioned earlier, these tools can be categorized into two types. Those that only allow simulation of an edge learning environment in the cloud, and those that allow training on edge devices.
Simulation only FL tools
@@
Simulation tools are extremely important in the context of edge computing. They are used to model the behavior of fog/edge infrastructures, allowing for the study of interoperability across different layers and protocols in edge-cloud environments [382]. In edge learning, simulators enable experimentation with ML models on cloud servers, facilitating rapid prototyping and experimentation. Many edge learning simulators focus on FL. Notable examples include FL_PyTorch [377] a PyTorch-based simulation tool for FL, and TensorFlow Federated [^10] https://www.tensorflow.org/federated a similar tool for TensorFlow. Other notable FL simulators are: LEAF [372]; FedJax [371] a JAX-based open source library; Flute [373] an open source platform with multiple optimization, privacy, and communication strategies; And finally FedLab [370] a lightweight open-source framework that focus on algorithm effectiveness and communication efficiency, and allows customization on server optimization, client optimization, communication agreement, and communication compression.
FL Simulation tools, which allow the training on an edge device
In recent years, numerous libraries have emerged to support the experimentation, development, and deployment of FL algorithms on edge devices. These libraries enable researchers to develop and test FL algorithms on the cloud while facilitating the transition from simulation to real-world deployment on the edge. Notable examples include Flower [366], and FedML [367] which are both aimed toward the research and experimentation of FL algorithms, while allowing the execution of the algorithms on a variety of edge devices. PySyft [368] is another FL open-source library that was built as an extension of PyTorch, Keras, and TensorFlow, and can be run on mobile devices using KotlinSyft [^11] KotlinSyft: Syft worker for secure on-device machine learning for Android https://github.com/OpenMined/KotlinSyft for Android and SwiftSyft [^12] SwiftSyft: Syft worker for secure on-device machine learning for iOS https://github.com/OpenMined/SwiftSyft for iOS. FedERA [369] is a similar library that includes a verification module to ensure the validation of local models and avoid aggregating malicious ones. Additionally, FedERA features a carbon emission tracker module to accurately estimate CO2 emissions during the local parameter update phase.
Other non-FL Frameworks for edge learning
Edge learning frameworks and libraries have been proposed to target specific platforms or hardware. These frameworks aim to simplify the training of ML models on various devices. Notable examples include CoreML, a Swift-based ML inference and training framework for iOS, designed to simplify ML model deployment and training on iOS devices. PULP-TrainLib [379] is another framework, proposed for on-device training on RISC-V multi-core microcontrollers. Additionally, EdgeRL [378] is a lightweight C/C++ framework for on-device reinforcement learning, designed to run on single-core processors typically found in resource-limited embedded platforms.
VII Open issues, research directions and future trends
In this section, we dive into the challenges, emerging research paths, and future trends in edge learning. To provide a comprehensive overview, we will divide this section into two parts. The first part will examine the open issues and existing challenges in edge learning, highlighting the obstacles that need to be addressed. The second part will explore promising research directions and our predictions for future trends, shedding light on the opportunities and possibilities that lie ahead.
VII-A Challenges and open issues
VII-A1 Resource constraints
As highlighted in previous parts of this survey, the resource constraints inherent to edge devices pose the biggest challenge for edge learning. In Section III we explored the different approaches designed to optimize and accelerate the training of ML models on the edge. However, despite current efforts, limitations in computation, memory, and sometimes energy continue to impede the training of the largest and most complex ML models on the edge. As recent tasks and use cases demand bigger and more complex ML models, the resource limitations of edge devices still pose a significant challenge and remain an open issue. Consequently, ongoing research efforts focus on optimizing ML models for resource constrained environments. Another promising idea, not explored in this survey, involves optimizing edge device hardware for ML [383, 384, 385].
VII-A2 Challenges in detecting data quality issues in the edge
Ensuring data quality is a crucial aspect of training ML models [386, 387]. However, this has proven challenging in the context of edge learning. Due to the decentralized nature of storage inherent in edge computing, detecting data quality issues such as missing or incorrect labels and noisy data is difficult. Additionally, edge devices can be prone to hardware failures, leading to missing or corrupted data [388, 389]. As such, data quality for ML on the edge remains an ongoing challenge [390], necessitating further research into developing new methods for detecting and fixing data quality issues, as well as designing ML models that are robust to these issues. The survey [390] explores the different challenges, constraints, potential solutions, and ongoing efforts related to data quality for ML on the edge.
VII-A3 Lack of labelled data availability